# Policy Learning with Confidence

### v1

@@ 2.1. Motivation and Risk-Aware Policy Choice
Hence, for each fund allocation $\pi$, the agency is given $\widehat{V}(\pi)$ along with its associated risk $s(\pi)$. What should the agency do? Our proposal is again to use the risk-offset rule (4). If risk is ignored, one might rely on the empirical welfare maximizer, but in the absence of additional constraints, that rule would simply allocate all funds to the program with the highest estimated return. This is generally unappealing on intuitive grounds. In contrast, the risk-aware approach (4) would spread out funds in an “efficient” portfolio of programs, much like Markowitz portfolio allocation in finance.
2.2. Decisions Aware of Estimation Risk
@@
We consider a decision-maker (DM) who holds posterior beliefs $B$ about policy values $\{V(\pi)\}$ given by
@@
stemming from using the policy rule $\pi$ rather than the unknown welfare-maximizing choice, and the estimation risk component $\bigl{|}\widehat{V}(\pi)-V(\pi)\bigr{|}$. Consider the DM who minimizes the expected regret risk-aware loss function:
@@
Propositions 1 and 2 explain why DMs may prefer a strictly positive $k$, but the exact value of $k$ is application-specific. One way to set it is by conducting experiments where DMs respond to hypothetical or incentivized lotteries/choices, then calibrating $k$ based on their observed choices. Another approach is to present a range of efficient decision frontiers, have DMs select their preferred allocations, and reverse-engineer $k$ that aligns with those selections. Generally, there is no single universal way to determine the “right” $k$; its choice depends on context and stakeholder preferences. With that in mind, we first develop our theoretical results for general risk-aware decisions.
2.3. Regret Properties and Guarantees of Any Risk-Aware Policy
We now provide regret bounds for any risk-aware decision rule $\widehat{\pi}_{RW}(\widehat{k})$ that solves the risk-adjusted empirical welfare problem:
$\widehat{\pi}_{RW}(\widehat{k})\;\in\;\arg\max_{\pi\in\Pi}\bigl{\{}\widehat{V}(\pi)\;-\;\widehat{k}\,\widehat{s}(\pi)\bigr{\}},$
Scaled Estimation Error and Gaussian Approximation. Define
$\widehat{Z}_{\pi}\;:=\;\frac{\widehat{V}(\pi)-V(\pi)}{\widehat{s}(\pi)}$
  • (G)
These conditions are known to hold under mild assumptions. [^13] See Chernozhukov et al. (2022, 2023a) for sharp forms of Gaussian approximations, and Chernozhukov et al. (2023b) for a review. Additional references, e.g. Belloni et al. (2018); Quintas-Martinez (2022) discuss Gaussian approximations for many or continuum of target parameters, including those learned via debiased machine learning. Furthermore,  (G) implicitly requires consistency of risk estimates: $\max_{\pi\in\Pi}\lvert\widehat{s}(\pi)/s(\pi)-1\rvert\to_{P}0$.
denote the $(1-\beta)$-quantile of the estimation error over $K$ under the Gaussian approximation. Under (G), with probability at least $1-\beta-r_{n}$, $\max_{\pi\in K}\widehat{Z}_{\pi}\leq q_{1-\beta,K}$. Let $\Pi_{0}=\{\pi\in\Pi:V(\pi)=V_{\max}\}$ be the set of policies with maximal true value, and denote its cardinality by $p_{0}=|\Pi_{0}|$. Define
These represent, respectively, an upper bound on the risk of all policies in $\Pi$ and a lower bound on the risk of the best policies $\Pi_{0}$.
Proposition 3 (Regret Bounds for Risk-Aware Decisions).
Under condition (G), the regret of any RW policy selection rule $\widehat{\pi}_{RW}(\widehat{k})$ is bounded with probability at least $1-2\beta-2r_{n}$ as:
$V_{\max}-V(\widehat{\pi}_{RW}(\widehat{k}))\;\;\leq\;\;\frac{\underline{\sigma}_{\Pi_{0}}}{\sqrt{n}}\Bigl{\{}q_{1-\beta,\Pi_{0}}+\widehat{k}\Bigr{\}}\;+\;\frac{\bar{\sigma}_{\Pi}}{\sqrt{n}}\Bigl{\{}\bigl{(}q_{1-\beta,\Pi}-\widehat{k}\bigr{)}_{+}\Bigr{\}}.$
$V_{\max}-V(\widehat{\pi}_{RW}(\widehat{k}))\;\;\lesssim\;\;\bigl{(}\bar{\sigma}_{\Pi}/\sqrt{n}\bigr{)}\,q_{1-\beta,\Pi}.$
$V_{\max}-V(\widehat{\pi}_{RW}(\widehat{k}))\;\;\lesssim\;\;\bigl{(}\underline{\sigma}_{\Pi_{0}}/\sqrt{n}\bigr{)}\,q_{1-\beta,\Pi}.$
$\bar{\sigma}_{\Pi}\,\sqrt{\frac{\log p}{n}},$
as $p\to\infty$ (and $n\to\infty$). This class includes the empirical welfare maximizing rule ($\widehat{k}=0$) and matches known optimal minimax rates when available (see Athey and Wager (2021)). Additionally, choosing $\widehat{k}>0$ sufficiently large can yield a regret bound of
$\underline{\sigma}_{\Pi_{0}}\,\sqrt{\frac{\log p}{n}},$
$\underline{\sigma}_{\Pi_{0}}<\bar{\sigma}_{\Pi}$
if there is significant heterogeneity in policy risks.
Reporting Guarantees of Risk-Aware Decisions
  • (B)
    With probability at least $1-\delta_{n}$ (where $\delta_{n}\searrow 0$),
Proposition 4 (LCB Guarantees for RW Decisions).
2.4. The PoLeCe and Its Properties
We now introduce policy learning with confidence (PoLeCe), a risk-aware rule that provides a key reporting guarantee by maximizing the LCB on policy value (welfare):
Notably, PoLeCe is an RW decision that uses
$\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$
serves as a natural performance guarantee.
Remark 1 (PoLeCe as Minimizer of the UCB on Regret).
Proposition 5 (Reporting Guarantees and Regret Bounds for PoLeCe).
With probability at least $1-\alpha-r_{n}^{\prime}$,
$V_{\max}\;\;\geq\;\;V(\widehat{\pi}_{\mathrm{PoLeCe}})\;\;\geq\;\;LV_{1-\alpha,\Pi}.$
where $\underline{\sigma}_{\Pi_{0}}=\sqrt{n}\,\min_{\pi\in\Pi_{0}}\widehat{s}(\pi)$ is determined by the least risky policy among those with maximal true policy value.
Here, $LV_{1-\alpha,\Pi}$ provides a valid high-confidence lower bound on the true policy value of the PoLeCe choice. The second result in Proposition 5 refines the general regret bounds from Proposition 3. As noted after Proposition 3, $\beta$ captures the trade-off between failure probability and bound tightness. The constant $\alpha$ determines the confidence level of the lower bound $LV_{1-\alpha,\Pi}$ and is chosen by the researcher. When $\alpha$ is sufficiently small, Proposition 5 provides informative bounds. However, for larger $\alpha$ (e.g., $\alpha=0.5$), the bounds above become uninformative, in which case the general bounds from Proposition 3 remain useful, as they apply to all risk-aware decision rules.
Another key observation is that the leading constant in the regret bound is governed by the least risky among the best policies, and can therefore be significantly smaller than that for other risk-aware rules (including EWM) in settings with substantial risk heterogeneity. This theoretical advantage is corroborated by the computational experiments we present below.
@@ 2.5.2. Discretization for Theoretical Analysis
If the original set of policies, denoted by $\Pi^{o}$, is infinite, we do replace it with a finite $\Pi$ of cardinality $p$. We need a discretization that is both fine enough and not overly complex, so that with probability at least $1-\delta_{n}$,
@@
2.6. Computational Experiments
Table 1. Calibrated Simulation Results
Average Median 95%-percentile Average
regret regret regret Welfare LCB
Measles shots EWM 7.88% 2.12% 32.10% -36.20%
(DGP 1) PoLeCe 8.83% 10.02% 12.84% -21.53%
Control 52.64% 52.64% 52.64% -79.92%
Shots per dollar EWM 1.74% 1.77% 4.89% -6.64%
(DGP 2) PoLeCe 1.00% 0.00% 3.83% -5.98%
Control 4.89% 4.89% 4.89% -16.31%
Informal Saving Technology EWM 15.91% 17.21% 33.85% -54.17%
(DGP 3) PoLeCe 28.14% 30.77% 41.45% -45.02%
Control 66.39% 66.39% 66.39% -87.82%
Alcohol and Self-Control EWM 1.38% 1.08% 4.46% -3.66%
(DGP 4) PoLeCe 0.84% 0.47% 2.21% -3.46%
Control 1.43% 1.43% 1.43% -7.86%
@@ 3.1. Investment Allocations for Public Programs
This section continues the discussion from the introduction, and illustrates how to solve the problem of investment allocations for public programs described in Section 2.1. We obtain MVPF estimates $\widehat{R}$ from the Policy Impacts Library, Hendren and Sprung-Keyser (2021). There are 172 programs in total, and we focus on those in the United States whose MVPF is estimated by an RCT. We assume the reported upper and lower bounds correspond to 95% confidence intervals and infer the standard errors for each $\widehat{R}$ from these bounds. [^16] We therefore exclude programs for which the upper and lower bounds are infinite. This yields $14$ programs. We present their MVPF estimates and standard errors in Table 2 and provide the Policy Impacts Library’s description of each corresponding social program in Appendix D. We also assume the estimation errors across programs are independent, so their covariance matrix is diagonal, with the squared standard errors on the diagonal. Note that here we only account for estimation error, holding the specification in the original papers that produced these MVPF estimates fixed. Alternative specifications can lead to different MVPF estimates for the same program. For example, accounting for transfers to parents could lead to larger MVPF estimates for early childhood programs.
Table 2. MVPF estimates and standard errors for selected programs.
Program MVPF S.E.
Holistic Wrap-around Services Can Improve Employment Rates 1.39 1.11
Padua Pilot 0.18 0.23
Head Start Impact Study 1.84 0.47
Wisconsin Scholar Grant to Low-Income College Students 1.43 0.38
Housing Vouchers in Chicago 0.65 0.02
Oregon Health Insurance Experiment (Provided to Single Adults) 1.16 0.04
Income Maintenance Experiment in Seattle and Denver 0.27 0.95
Paycheck Plus: EITC to Adults without Dependents 1.00 0.09
Work Advance 0.78 0.28
JobStart 0.20 0.11
Year Up 0.43 0.03
Job Corps 0.15 0.22
Job Training Partnership Act, Adults 1.38 0.40
Job Training Partnership Act, Youth -0.23 0.81
@@
Table 3 reports the allocation selected by PoLeCe and EWM if we restrict the set of programs to those targeting individuals older than 25. As expected, the EWM rule is a corner solution that places all investment in the best-performing program according to the empirical estimates, Holistic Wrap-around Services Can Improve Employment Rates, as its MVPF estimate is the highest. The PoLeCe method, in sharp contrast to EWM, places a large weight on housing vouchers in Chicago and Medicaid for single adults as their MVPF estimates are among the highest, and are highly precise. The remaining share is distributed across programs that have high MVPF estimates but also larger standard errors, which reflects the high potential benefit of these programs while accounting for the statistical imprecision of their MVPF estimates.
@@
Esimated Value: $\widehat{V}(\widehat{\pi})$ 1.39 1.17
@@
Panel B. Optimal Treatment Policy
Holistic Wrap-around Services Can Improve Employment Rates 1 0.00
Padua Pilot 0 0.01
Housing Vouchers in Chicago 0 0.68
Oregon Health Insurance Experiment (Provided to Single Adults) 0 0.25
Income Maintenance Experiment in Seattle and Denver 0 0.00
Paycheck Plus: EITC to Adults without Dependents 0 0.05
Work Advance 0 0.01
Job Training Partnership Act, Adults 0 0.00
Note: EWM denotes empirical welfare maximization, and PoLeCe denotes policy learning with confidence. In Panel A, we report the EWM and the PoLeCe lower confidence bound with $\alpha=0.05$. In Panel B, we report the optimal treatment policy for each program.
@@
Panel B. Optimal Treatment Policy
Head Start Impact Study 1 0.52
Wisconsin Scholar Grant to Low-Income College Students 0 0.48
JobStart 0 0
Year Up 0 0
Job Corps 0 0
Job Training Partnership Act, Youth 0 0
Note: EWM denotes empirical welfare maximization, and PoLeCe denotes policy learning with confidence. In Panel A, we report the EWM and the PoLeCe lower confidence bound with $\alpha=0.05$. In Panel B, we report the optimal treatment policy for each program.

### v2

@@ 2.1. Motivation and Risk-Aware Policy Choice
Hence, for each fund allocation $\pi$, the agency is given $\widehat{V}(\pi)$ along with its associated estimation risk $s(\pi)$. What should the agency do? Our proposal is again to use the risk-aware rule (4). If estimation risk is ignored, one might rely on the empirical welfare maximizer, but in the absence of additional constraints, that rule would simply allocate all funds to the program with the highest estimated return. This is generally unappealing on intuitive grounds. In contrast, the risk-aware approach (4) would spread out funds in an “efficient” portfolio of programs, similar to Markowitz portfolio allocation in finance. However, the motivation and formulation of our approach are distinct from the Markowitz model.
2.2. Regret Properties and Guarantees of Any Risk-Aware Policy
We now provide regret bounds for any risk-aware decision rule $\widehat{\pi}_{\text{RW}}(\widehat{k})$ that solves the risk-adjusted empirical welfare problem:
$\widehat{\pi}_{\text{RW}}(\widehat{k})\;\in\;\arg\max_{\pi\in\Pi}\bigl{\{}\widehat{V}(\pi)\;-\;\widehat{k}\,\widehat{s}(\pi)\bigr{\}},$
Scaled Estimation Error and Gaussian Approximation. Define
$\widehat{Z}_{\pi}\;:=\;\frac{\widehat{V}(\pi)-V(\pi)}{\widehat{s}(\pi)}$
  • (G)
These conditions are known to hold under mild assumptions. [^11] See Chernozhukov et al. (2022, 2023a) for sharp forms of Gaussian approximations, and Chernozhukov et al. (2023b) for a review. Additional references, e.g. Belloni et al. (2018); Quintas-Martinez (2022) discuss Gaussian approximations for many or a continuum of target parameters, including those learned via debiased machine learning. Furthermore,  (G) implicitly requires consistency of risk estimates: $\max_{\pi\in\Pi}\lvert\widehat{s}(\pi)/s(\pi)-1\rvert\to_{P}0$.
denote the $(1-\beta)$-quantile of the estimation error over $K$ under the Gaussian approximation. Under (G), with probability at least $1-\beta-r_{n}$, $\max_{\pi\in K}\widehat{Z}_{\pi}\leq q_{1-\beta,K}$. Let $V_{\max}=\sup_{\pi\in\Pi}V(\pi)$ be the maximal true welfare and $\Pi_{0}=\{\pi\in\Pi:V(\pi)=V_{\max}\}$ be the set of policies with maximal true welfare, and denote its cardinality by $p_{0}=|\Pi_{0}|$. Define
These represent, respectively, an upper bound on the estimation risk of all policies in $\Pi$ and a lower bound on the estimation risk of the best policies $\Pi_{0}$.
Proposition 1 (Regret Bounds for Risk-Aware Decisions).
Under condition (G), the regret of any RW policy selection rule $\widehat{\pi}_{\text{RW}}(\widehat{k})$ is bounded with probability at least $1-2\beta-2r_{n}$ as:
$V_{\max}-V(\widehat{\pi}_{\text{RW}}(\widehat{k}))\;\;\leq\;\;\frac{\underline{\sigma}_{\Pi_{0}}}{\sqrt{n}}\Bigl{\{}q_{1-\beta,\Pi_{0}}+\widehat{k}\Bigr{\}}\;+\;\frac{\bar{\sigma}_{\Pi}}{\sqrt{n}}\Bigl{\{}\bigl{(}q_{1-\beta,\Pi}-\widehat{k}\bigr{)}_{+}\Bigr{\}}.$
$V_{\max}-V(\widehat{\pi}_{\text{RW}}(\widehat{k}))\;\;\lesssim\;\;\bigl{(}\bar{\sigma}_{\Pi}/\sqrt{n}\bigr{)}\,q_{1-\beta,\Pi}.$
$V_{\max}-V(\widehat{\pi}_{\text{RW}}(\widehat{k}))\;\;\lesssim\;\;\bigl{(}\underline{\sigma}_{\Pi_{0}}/\sqrt{n}\bigr{)}\,q_{1-\beta,\Pi}.$
$\bar{\sigma}_{\Pi}\,\sqrt{\frac{\log p}{n}},$
as $p\to\infty$ (and $n\to\infty$). This class includes the empirical welfare maximization rule ($k=0$) and matches known optimal minimax rates when available (see Athey and Wager (2021)). Additionally, choosing $\widehat{k}>0$ sufficiently large can yield a regret bound of
$\underline{\sigma}_{\Pi_{0}}\,\sqrt{\frac{\log p}{n}},$
$\underline{\sigma}_{\Pi_{0}}<\bar{\sigma}_{\Pi}$
if there is significant heterogeneity in estimation risks across policies. The empirical risk minimization literature has also used penalties on estimation risk to improve error bounds, albeit in different contexts; see, for example, Maurer and Pontil (2009); Swaminathan and Joachims (2015); Foster and Syrgkanis (2023).
Reporting Guarantees of Risk-Aware Decisions
  • (B)
    With probability at least $1-\delta_{n}$ (where $\delta_{n}\searrow 0$),
Proposition 2 (LCB Guarantees for RW Decisions).
2.3. The PoLeCe Rule and Its Properties
We now introduce policy learning with confidence (PoLeCe), a risk-aware rule that provides a key reporting guarantee by maximizing the LCB on policy welfare:
Notably, PoLeCe is an RW decision that uses
$\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$
serves as a natural performance guarantee.
Remark 1 (PoLeCe as Minimizer of the UCB on Regret).
Proposition 3 (Reporting Guarantees and Regret Bounds for PoLeCe).
With probability at least $1-\alpha-r_{n}^{\prime}$,
$V_{\max}\;\;\geq\;\;V(\widehat{\pi}_{\mathrm{PoLeCe}})\;\;\geq\;\;LV_{1-\alpha,\Pi}.$
where $\underline{\sigma}_{\Pi_{0}}=\sqrt{n}\,\min_{\pi\in\Pi_{0}}\widehat{s}(\pi)$ is determined by the least risky policy among those with maximal true policy welfare.
The first result in Proposition 3 shows that $LV_{1-\alpha,\Pi}$ provides a valid high-confidence lower bound on the true policy welfare of the PoLeCe choice. The second result sharpens the general regret bounds presented in Proposition 1. As noted following Proposition 1, the parameter $\beta$ governs the trade-off between failure probability and bound tightness, while the constant $\alpha$, chosen by the researcher, determines the confidence level of the lower bound $LV_{1-\alpha,\Pi}$. The refinement in Proposition 3 depends on the relationship between $\alpha$ and $\beta$. When $\alpha=\beta$, Propositions 1 and 3 yield qualitatively identical guarantees. When $\alpha<\beta$, the bounds in both propositions remain qualitatively identical, but the bound in Proposition 3 holds with higher probability. When $\alpha>\beta$, the bound in Proposition 3 is tighter, but it may hold with lower probability. For a large $\alpha$, the bound may fail to hold; in such cases, the general bound from Proposition 1 remains valid and informative, as it applies to all risk-aware decision rules.
Another key observation is that the leading constant in the regret bound is governed by the least risky among the best policies, and can therefore be significantly smaller than that for the EWM rule in settings with substantial risk heterogeneity. This theoretical advantage is corroborated by the computational experiments we present below.
2.4. Computational Experiments
Table 1. Calibrated Simulation Results
Average Median 95%-percentile Average
regret regret regret Welfare LCB
Measles shots EWM 7.88% 2.12% 32.10% -36.20%
(DGP 1) PoLeCe 8.83% 10.02% 12.84% -21.53%
Control 52.64% 52.64% 52.64% -79.92%
Shots per dollar EWM 1.74% 1.77% 4.89% -6.64%
(DGP 2) PoLeCe 1.00% 0.00% 3.83% -5.98%
Control 4.89% 4.89% 4.89% -16.31%
Informal Saving Technology EWM 15.91% 17.21% 33.85% -54.17%
(DGP 3) PoLeCe 28.14% 30.77% 41.45% -45.02%
Control 66.39% 66.39% 66.39% -87.82%
Alcohol and Self-Control EWM 1.38% 1.08% 4.46% -3.66%
(DGP 4) PoLeCe 0.84% 0.47% 2.21% -3.46%
Control 1.43% 1.43% 1.43% -7.86%
2.5. Alternative Motivations for Risk-Aware Decisions
@@
We consider a decision-maker (DM) who holds posterior beliefs $B$ about policy welfare $\{V(\pi)\}$ given by
@@
stemming from using the policy rule $\pi$ rather than the unknown welfare-maximizing choice, and the estimation risk component $\bigl{|}\widehat{V}(\pi)-V(\pi)\bigr{|}$. Consider a DM who minimizes the expected regret risk-aware loss function:
@@
Propositions 4 and 5 explain why DMs may prefer a strictly positive $k$, but the exact value of $k$ is application-specific. One way to set it is by conducting experiments where DMs respond to hypothetical or incentivized lotteries/choices, then calibrating $k$ based on their observed choices. Another approach is to present a range of efficient decision frontiers, have DMs select their preferred allocations, and reverse-engineer $k$ that aligns with those selections. Generally, there is no single universal way to determine the “right” $k$; its choice depends on context and stakeholder preferences. Using a critical value for $k$ as in PoLeCe provides both interpretable regret and reporting guarantees.
@@ 2.6.2. Discretization for Theoretical Analysis
If the original set of policies, denoted by $\Pi^{o}$, is infinite, we replace it with a finite $\Pi$ of cardinality $p$. In this case we need a discretization that is both fine enough and not overly complex, so that with probability at least $1-\delta_{n}$,
@@ 3.1. Investment Allocations for Public Programs
This section continues the discussion from the introduction, and illustrates how to solve the problem of investment allocations for public programs described in Section 2.1. We obtain MVPF estimates $\widehat{R}$ from the Policy Impacts Library, Hendren and Sprung-Keyser (2021). There are 172 programs in total, and we focus on those in the United States whose MVPF is estimated by an RCT. We assume the reported upper and lower bounds correspond to 95% confidence intervals and infer the standard errors for each $\widehat{R}$ from these bounds. [^16] We therefore exclude programs for which the upper and lower bounds are infinite. This yields $14$ programs and we provide the Policy Impacts Library’s description of each corresponding social program in Appendix D. We also assume the estimation errors across programs are independent, so their covariance matrix is diagonal, with the squared standard errors on the diagonal. Note that here we only account for estimation error, holding the specification in the original papers that produced these MVPF estimates fixed. Alternative specifications can lead to different MVPF estimates for the same program. For example, accounting for transfers to parents could lead to larger MVPF estimates for early childhood programs.
@@
Table 2 reports the allocation selected by PoLeCe and EWM if we restrict the set of programs to those targeting individuals older than 25. As expected, the EWM rule is a corner solution that places all investment in the best-performing program according to the empirical estimates, Holistic Wrap-around Services Can Improve Employment Rates, as its MVPF estimate is the highest. The PoLeCe method, in sharp contrast to EWM, places a large weight on housing vouchers in Chicago and Medicaid for single adults as their MVPF estimates are among the highest, and are highly precise.
@@
Estimated Value: $\widehat{V}(\widehat{\pi})$ 1.39 1.17
@@
Panel B. Selected Policy
Holistic Wrap-around Services Can Improve Employment Rates: 1.39 (1.11) 1 0.00
Padua Pilot: 0.18 (0.23) 0 0.01
Housing Vouchers in Chicago: 0.65 (0.02) 0 0.68
Oregon Health Insurance Experiment (Provided to Single Adults): 1.16 (0.04) 0 0.25
Income Maintenance Experiment in Seattle and Denver: 0.27 (0.95) 0 0.00
Paycheck Plus: EITC to Adults without Dependents: 1.00 (0.09) 0 0.05
Work Advance: 0.78 (0.28) 0 0.01
Job Training Partnership Act, Adults: 1.38 (0.40) 0 0.00
Note: EWM denotes empirical welfare maximization, and PoLeCe denotes policy learning with confidence. In Panel A, we report the EWM and the PoLeCe lower confidence bound with $\alpha=0.05$. In Panel B, we report the selected allocation for each program. MVPF estimates and associated standard errors (in parentheses) are included alongside each program name.
@@
Panel B. Selected Policy
Head Start Impact Study: 1.84 (0.47) 1 0.52
Wisconsin Scholar Grant to Low-Income College Students: 1.43 (0.38) 0 0.48
JobStart: 0.20 (0.11) 0 0
Year Up: 0.43 (0.03) 0 0
Job Corps: 0.15 (0.22) 0 0
Job Training Partnership Act, Youth: -0.23 (0.81) 0 0
Note: EWM denotes empirical welfare maximization, and PoLeCe denotes policy learning with confidence. In Panel A, we report the EWM and the PoLeCe lower confidence bound with $\alpha=0.05$. In Panel B, we report the selected allocation for each program. MVPF estimates and associated standard errors (in parentheses) are included alongside each program name.